OpenStudy (anonymous):

a customer buys 4 hot dogs and 5 bags of potato chips for $12 another customer buys 3 hot dog and 4 chips for $9.25. find the cost of each item ,

OpenStudy (anonymous):

let x=number of hot dogs, y=number of bags of potato chips. 4x+5y=12, 3x+4y=9.25. solve these equations simultaneously :)

OpenStudy (anonymous):

lets say one hot dog costs x dollars lets say one bag of potato chips costs y dollars. 4 hot dogs cost 4x dollars. 5 bags of chips cost 5y dollars. 4 hot dogs and 5 bags of chips together cost 12 dollars. so, 4x+5y = 12. similarly get the other part in the form of an equation too and solve it for x and y

OpenStudy (radar):

better assignment of the variables. Letting x = cost, rather than number of dogs and chips. That is already known.

OpenStudy (anonymous):

the answer was given each hot dog cause $1.75 and each bag of chips cause $1 each i want to know ow there got this answer

OpenStudy (anonymous):

how they got that answer

OpenStudy (anonymous):

It has already been explained to you. do you know how to solve two simultaneous equations?

OpenStudy (anonymous):

my teacher want us to explain how he got that answer

OpenStudy (anonymous):

no i dont

OpenStudy (anonymous):

so you dont know how to solve x+y = 5 x-y = 1 for x and y?

OpenStudy (anonymous):

i do,

OpenStudy (anonymous):

then what is the problem? those are simultaneous equations.

OpenStudy (anonymous):

i just want to know how he got $1.75 for the hot dog and $1 for the chips

OpenStudy (anonymous):

refer to what I said earlier. lets say one hot dog costs x dollars. how much will 4 hot dogs cost?

OpenStudy (anonymous):

4

OpenStudy (anonymous):

please help me dont get frustrated with me, i

OpenStudy (anonymous):

4 dollars

OpenStudy (anonymous):

say one hot dog costs 2 dollars. how much will 4 hot dogs cost?

OpenStudy (anonymous):

times 2 by 4

OpenStudy (anonymous):

ok good. so if one hot dog costs x dollars, how much will 4 hot dogs cost?

OpenStudy (anonymous):

a $ 1

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

sorry

OpenStudy (anonymous):

i am so confuse

OpenStudy (anonymous):

if u can please explain to me then we find the solution later i will be so happy, is soon due

OpenStudy (anonymous):

answer the question. you dont know how much each hot dog costs. So lets say we assign the variable x to the cost of each hot dog. so one hot dog costs x amount of dollars. suppose you went to the store and purchased one hot dog for x amount of dollars, how much would you have to pay to buy 2 hot dogs?

OpenStudy (anonymous):

x dollar

OpenStudy (anonymous):

are you sure? that is the cost of one hot dog. what is the cost of 2 hot dogs?

OpenStudy (anonymous):

you gave x dollars to purchase one hot dog. you have to give another x dollars to purchase one more hot dog. Is that correct or not?

OpenStudy (anonymous):

is correct

OpenStudy (anonymous):

okay so to purchase 2 hot dogs, you gave x dollars + another x dollars. so that is 2 times x dollars. Is that correct or not?

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

so how many x dollars do you need to give to purchase 4 hot dogs?

OpenStudy (anonymous):

2x

OpenStudy (anonymous):

oh sorry 4X

OpenStudy (anonymous):

okay good. so to purchase 4 hot dogs you have to give 4x dollars. similarly, say one bag of chips costs y dollars. how much does it cost to purchase 5 bags of chips?

OpenStudy (anonymous):

5dollars

OpenStudy (anonymous):

can you please help me with another problem

OpenStudy (anonymous):

didn't I just explain this to you? you purchase one bag of chips for y dollars. you have to give another y dollars to purchase one more bag of chips. so how many dollars do you have to give to purchase 5 bags?

OpenStudy (anonymous):

thanks for this one, i have another i need help

OpenStudy (anonymous):

I have a good piece of advice for you. Do not take math as a major course.

OpenStudy (anonymous):

a university theater sold 556 tickets for a play, tickets cost $22 per adult tickets and $ 12 for senior citizen ticket, if the total receipt was 8492, how many senior citizen ticket were sold

OpenStudy (anonymous):

i know rite, i am just talking this because i need it to graduate, i dont like math, i am more a science student

OpenStudy (anonymous):

I hope you are not a physics student then :)

OpenStudy (anonymous):

let say there are a certain number of adult tickets and a certain number of senior citizen tickets. let us say there are A number of adult tickets and S number of senior citizen tickets.

OpenStudy (anonymous):

total number of tickets is 556. so A+S =556. each adult ticket costs 22 dollars. so A number of adult tickets cost 22 times A amount of dollars. each senior citizen ticket costs 12 dollars. so S number of senior citizen tickets cost 12 times S amount of dollars. so, 22A+12S =8492. so you have two equations. A+S =556 22A +12S =8492. solve to get S and A

OpenStudy (anonymous):

i am a nursing student

OpenStudy (anonymous):

so you time 22*12?

OpenStudy (anonymous):

why would you do that?

OpenStudy (anonymous):

i was trying to use the multiplication sign

OpenStudy (anonymous):

I thought you knew how to solve simultaneous equations

OpenStudy (anonymous):

oh.. okay. there is a button for equations here: | | v

OpenStudy (anonymous):

because you said 22A PLUS 12S = 8492

OpenStudy (anonymous):

yes, thats right. but why are you multiplying 12 and 22?

OpenStudy (anonymous):

http://www.youtube.com/watch?v=XM7Q4Oj5OTc go through that video and all the other videos in the related videos section. there are many examples

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

so when we add A+S IT WILL BE 34, ARE WE DIVIDING IT WITH ANYTHING

OpenStudy (anonymous):

why is A+S = 34? don't you remember what we assigned the variable A to? A is the NUMBER of adult tickets, NOT the price of each adult ticket. we already know the price is 22.

OpenStudy (anonymous):

READ what I posted PROPERLY instead of asking these silly questions. This is starting to get frustrating for me. I have been patiently explaining all this to you.

OpenStudy (anonymous):

SORRY AND THANKS FOR UR HELP U HAVE BEEN REALLY PATIANT WITH ME TODAY I REALLY APPRACIATE IT

OpenStudy (anonymous):

you are welcome. :) think about these things logically. put yourself in the position of the problem and see what you would do in that situation. what if you were the manager of the theater and you had 556 tickets? what if you made 8492 dollars. How many adult tickets did you sell at 22 dollars per ticket and so on.

OpenStudy (anonymous):

THIS IS WHAT I CAME OUT WITH LET ME SAY X FOR ADULT Y FOR S.CITIZEN

OpenStudy (anonymous):

X+Y= 556 RIGHT

OpenStudy (anonymous):

I saw radar already gave you the answer. save yourself the trouble.

OpenStudy (anonymous):

I JUST WANTED TO KNOW IF I AM RIGHT RIGHT

OpenStudy (anonymous):

okay, post what you got then

OpenStudy (anonymous):

I GOT 372

OpenStudy (anonymous):

372 for what?

OpenStudy (anonymous):

372 ADULT TICKETS SOLD

OpenStudy (anonymous):

372 is wrong. it should be 374

OpenStudy (anonymous):

OH ARE U SERIOUS

OpenStudy (anonymous):

NOW I KNOW WHERE I MADE THE MISTAKE

OpenStudy (anonymous):

THANKS FOR PIONTING THAT OUT FOR ME

OpenStudy (anonymous):

no problem. http://www.analyzemath.com/Calculators/Calculator_syst_eq.html